Nico is saving money for his college education. He invests some money at 8​%, and ​$1500 less than that amount at 4 %. The investments produced a total of ​$180 interest in 1 yr. How much did he invest at each​ rate?
Answer by addingup(3677)

0.08x + 0.04(x-1500) = 180
0.08x + 0.04x - 60 = 180
0.12x = 240
x = 2000
He invested 2000 at 8% and 2000-1500 = 500 at 4%
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Check:
2000*0.08 = 160
500*0.04 = 20
. . . . . ---------
Totas . . . 180 Correct